OpenStudy (anonymous):

(x^2-2x+1)(y^2+2y+1)=49 factor

myininaya (myininaya):

is this written correctly? are we solving? we have two variables?

OpenStudy (anonymous):

factor the x then factor the y

myininaya (myininaya):

x^2-2x+1=(x-1)(x-1)=(x-1)^2 y^2+2y+1=(y+1)(y+1)=(y+1)^2

myininaya (myininaya):

this would be a circle if there was a plus in between (x^2-2x+1) and (y^2+2y+1)